OpenStudy (anonymous):

Find the distance between the points (2, -3) and (5, -4).

OpenStudy (whpalmer4):

The distance between two points \((x_1,y_1),(x_2,y_2)\) is given by the formula \[d = \sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\]Plug in your points and find the result. Notice that if the two points share either their x value or their y value, the formula reduces to the difference between the other values.

OpenStudy (whpalmer4):

It doesn't matter which point you choose as \((x_1,y_1)\)...

OpenStudy (anonymous):

d=(2-5)^2+(-3)-(-4)^2

OpenStudy (whpalmer4):

Close, \[d = \sqrt{(2-5)^2+(-3-(-4))^2}\]What does that equal?

OpenStudy (anonymous):

d=-23+169?

OpenStudy (whpalmer4):

Mmm...no :-) \[d = \sqrt{(2-5)^2 + (-3-(-4))^2}\]What is \((2-5)^2\)?

OpenStudy (anonymous):

-9?

OpenStudy (whpalmer4):

2-5= -3 -3*-3 = ?

OpenStudy (anonymous):

9?

OpenStudy (whpalmer4):

Right! Now what is -3 - (-4)?

OpenStudy (anonymous):

1?

OpenStudy (whpalmer4):

Correct again! So what is the square root of 9+1? That's our distance...

OpenStudy (whpalmer4):

You can't factor out any squares from 10, and it isn't a rational number, so you're left writing \[d = \sqrt{10}\]unless you want to break out your calculator and write down a few digits as an approximation.